HUBBLE SPACE TELESCOPE DAILY REPORT #1519 PERIOD COVERED: 0000Z (UTC) 11/16/95 - 0000Z (UTC) 11/17/95 Daily Status Report as of 321/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5957 (Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make observations for radio galaxy 3C206.0 = to study the morphology of that radio galaxy. The observations were = executed as scheduled, and no anomalies were noted. 1.2 Completed FOS/RD 5824 (Occultation of GSC5249 by Saturn) The Faint Object Spectrograph (Red Side) was used to observe = Saturn's atmosphere and its rings while being occulted by GSC5249. = Spectra of the occulted star, of Saturn's atmosphere and of Saturn's rings = will be obtained before and after the occultation. The observations were = completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flats) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.4 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 6303 (Observations of Galaxies Underlying Quasars) The WF/PC-2 was used to make observations of low red shift quasar = GSC5249-01 in order to study its underlying galaxy. The purpose is to = detect structure down to the 0.1 arcsecond level. The observations were = completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 6228 (Young Globular Clusters in Interacting and = Cooling Flow Galaxies) =3D=3D The WF/PC-2 was used to make observations of galaxy cluster = NGC-1275 looking for young globular clusters. Very deep exposures were = made of the target to attempt to distinguish between its two possible = formation histories. The observations were executed as scheduled, and no = anomalies were reported. 1.7 Completed WF/PC-2 6137 (A Search for Brown Dwarfs and Luminous Young = Planets in the Hyades) The WF/PC-2 was used to make observations of M class star RHY403 = which is located in the Hyades cluster, looking for planetary and brown = dwarf companions. The observations were executed as scheduled, and no = anomalies were noted. 1.8 Completed FGS 6051 (Dynamical Masses For The Stars in the Pre-Main = Sequence Spectroscopic Class) The FGSs were used to make astrometric measurements of the location = of star STAR-04560 to measure its relative orbit in a pre-main sequence = binary system. The observations were executed as scheduled, and no = anomalies were noted. 1.9 Completed WF/PC-2 5932 (Morphologies of High Redshift Radio Galaxies) The WF/PC-2 was used to make observations the high redshift galaxy = GAL-1140-2 (at distance Z>2) in order to study its morphologies, = especially the fine detail in clumping. The images were obtained in the = continuum and in Lyman-Alpha. During this proposal and as recorded in = HSTAR 5365, there was an FGS loss of lock. Otherwise, the observations = were executed as planned, and no further anomalies were noted. 1.10 Completed WF/PC-2 5454 (Ellipticals with Kinematically Distinct = Cores: Cycle 4 Medium) The WF/PC-2 was used to make observations of the small elliptical = galaxy NGC-4552 which contains a massive core which is kinematically = distinct from the rest of the galaxy. The observations were executed as = scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 6 Successful : 6 There was a loss of lock, occurring at 320/2043Z, that is reported in HSTAR 5365. 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: The Battery #2 Capacity Check began at 320/1209Z. The battery = reached 15.0 volts at 321/083228Z and was placed back on line at = 321/085640Z. An on-orbit PDB 43.B/PRS 6.5 test was conducted from 321/0300Z to = 321/1100Z. No problems were encoutered. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH